OpenStudy (anonymous):

A train traveled 300 miles. How long did the trip take if the train was traveling at a rate of: Use the d=rt formula a). (x + 10)mph b). (x - 5)mph A train traveled 300 miles. How long did the trip take if the train was traveling at a rate of: Use the d=rt formula a). (x + 10)mph b). (x - 5)mph @Mathematics

OpenStudy (anonymous):

cross check ur question

OpenStudy (anonymous):

\[t=\frac{d}{r} \]\[\frac{300}{x+10} \text{ and } \frac{300}{x-5} \]

OpenStudy (anonymous):

@robtobey do i go further im just asking

OpenStudy (anonymous):

There is not enough information given. You still are left with two unknowns at this point, t and x. All that is given as numbers is the distance, 300 miles

OpenStudy (anonymous):

ok so tell me if i got this right ok d=r*t 70mph d=300*70 d=21000

OpenStudy (anonymous):

70mph is not mentioned in your problem statement that you wrote.